# Superallowed $0^{+}\rightarrow 0^{+}$ $\beta$ decay studies at GANIL and upcoming opportunities with DESIR and S3-LEB

### v1

@@ 1 Introduction
Precision measurements of observables in superallowed (SA) Fermi $\beta$ transitions offer excellent probes to hunt for signatures of physics beyond the Standard Model (BSM) [1]. These decays proceed from the $J^{\pi}=0^{+}$ ground-state of the parent nucleus to the $J^{\pi}=0^{+}$ isobaric analog state in the daughter. In these decays, only the vector part of the weak interaction contributes. Consequently, the hypothesis of a conserved vector current (CVC) [2] predicts that the transition rate ($ft$ value) remains constant across all SA decays for a given isospin, irrespective of the parent-daughter pair. This hypothesis has been tested and verified to a precision of $5\times 10^{-4}$ [3] once individual $ft$ values are corrected for nuclear structure and radiative effects (denoted $\mathcal{F}t$). The validation of CVC thus allows to derive $V_{ud}$, the dominant top-row element of the Cabibbo-Kobayashi-Maskawa (CKM) matrix, that describes the mixing of quark flavor states with the mass eigenstates. The unitarity of this CKM matrix is fundamental to the current version of the Standard Model and is tested to the highest precision using the first-row elements. Being the largest and the most precise among the three elements, precision on $V_{ud}$ is key to performing sensitive tests of CKM unitarity. In fact, the highest precision on $V_{ud}$ currently comes from $\mathcal{F}t$ values of superallowed $0^{+}\rightarrow 0^{+}$ beta decays ($\mathcal{F}t^{0^{+}\rightarrow 0^{+}}$[3, 4]. Until recently, CKM unitarity was respected, however, a recent global re-evaluation of the $\mathcal{F}t^{0^{+}\rightarrow 0^{+}}$ values has indicated a tension with unitarity at the $2\sigma$ level [3, 4].
@@
Observables in SA $0^{+}\rightarrow 0^{+}$ $\beta$ decays, namely branching ratio (BR), half-life ($t_{1/2}$), and Q-value, are used to determine the experimental $ft^{0^{+}\rightarrow 0^{+}}$ value for the decay [4]. Individual $ft^{0^{+}\rightarrow 0^{+}}$ values, shown in the left panel of Fig. 1, are corrected for nuclear medium dependent isospin symmetry breaking (ISB) and radiative effects [3] to obtain the nucleus independent $\mathcal{F}t^{0^{+}\rightarrow 0^{+}}$ values shown in the right panel of Fig. 1. The average of these $\mathcal{F}t^{0^{+}\rightarrow 0^{+}}$ values is then used to derive $V_{ud}$
$\mathcal{F}t^{0^{+}\rightarrow 0^{+}}=ft^{0^{+}\rightarrow 0^{+}}(1+\delta^{{}^{\prime}}_{R})\left(1+\delta_{NS}-\delta_{C}\right)=\frac{K}{2V_{ud}^{2}G_{F}^{2}\left(1+\Delta_{R}\right)}$ (1)
@@
Here, the right hand side forms a constant with $K~{}=~{}8120.2787(11)\texttimes 10^{-10}$ GeV-4s and the nucleus-independent radiative correction $\Delta_{R}$, which is the same for all $0^{+}\rightarrow 0^{+}$ SA $\beta$ transitions [5]. In the middle term of Eq. 1, $\delta^{{}^{\prime}}_{R}$ and $\delta_{NS}$ are radiative corrections, and $\delta_{C}$ is the isospin symmetry breaking correction. These corrections range from $\sim-0.5\%$ to $\sim 2\%$ and among the three, $\delta_{C}$ increases strongly with the atomic charge $Z$. This is natural, as the Coulomb interaction grows stronger with $Z$, however this also enhances isospin mixing and symmetry breaking effects, consequently affecting the average $\mathcal{F}t^{0^{+}\rightarrow 0^{+}}$. As the average of $\mathcal{F}t^{0^{+}\rightarrow 0^{+}}$ defines the absolute value of $V_{ud}$, it is absolutely necessary to constrain these correction terms.
@@
Currently, the $ft^{0^{+}\rightarrow 0^{+}}$ values are measured for 23 SA $\beta$ emitters, ranging from 10C $(Z=6)$ to 74Rb $(Z=37)$ [3]. However, only 15 of these, shown in Fig. 1, are used in the evaluation of $V_{ud}$ because their $ft^{0^{+}\rightarrow 0^{+}}$ values have the required precision of $0.3\%$ or better [3]. For the remaining eight cases, that are shown in the left panel of Fig. 2, the uncertainty is dominated by those on the BR and Q-value (used to determine $f$ in Fig 2). Furthermore, as one can observe in the right panel of Fig. 2, there are large variations between ISB corrections calculated using different nuclear structure models. However, at present, only the shell model calculation using the Wood-Saxon wave function (annotated as SM-WS in Fig. 2) is used to correct the $ft^{0^{+}\rightarrow 0^{+}}$ values for ISB effects. To effectively constrain these corrections, a precision of $0.3\%$ or better is required for the experimental $ft$ values [3].
Figure 2: Left panel: Uncertainty budget on the corrected $\mathcal{F}t^{0^{+}\rightarrow 0^{+}}$ for the eight less precise cases of superallowed $\beta$ emitters. Figure created with data from Ref. [3]. Right panel: ISB corrections from different models for all known $0^{+}\rightarrow 0^{+}$ SA decays. Figure adapted from Ref [6].
An example is shown in Fig. 3 for the case of the SA emitter 18Ne, where a precision of $0.3\%$ on the branching ratio can help distinguish between two advanced models. Given the current tension with CKM unitarity, it is crucial to test these ISB models against stringent constraints set by experimental data, especially for the heavier superallowed emitters, where $\delta_{C}$ is most sensitive.
Figure 3: Top panel: $\mathcal{F}t^{0^{+}\rightarrow 0^{+}}$ obtained using two different nuclear structure models that calculate the ISB corrections for $ft^{0^{+}\rightarrow 0^{+}}$. Black data points correspond to shell model based corrections from Ref [3], referred to as SM-WS in Fig. 2. Red data points correspond to density functional based corrections of Ref [7], referred to as SV-DFT in Fig. 2. The two straight lines represent the average $\mathcal{F}t^{0^{+}\rightarrow 0^{+}}$ calculated using the 15 cases shown in Fig. 2. The inset displays the $\mathcal{F}t^{0^{+}\rightarrow 0^{+}}$ value for 18Ne dtermined using existing experimental data and the ISB corrections of Ref [3].
Bottom panel: $\delta_{C}$ corrections from Ref [3] and [7]. For $\delta_{C}=1$, using the CVC filter [3], the purple error bar demonstrates the potential to distinguish between the two models if the precision on the BR for 18Ne is 0.3%.
@@ 2 Experimental program
At GANIL, an experimental program is currently underway aiming towards constraining ISB corrections by performing high-precision measurements of the half-life and branching ratio of the 8 less-precisely known SA $\beta$ emitters. Of these eight nuclei, the first six nuclei (shown in Fig. 2) are $T_{Z}=-1$ parents, where the $\delta_{C}$ corrections are larger than those in $\beta$ decay pairs with $T_{Z}=0$ parents of the same mass.
@@
As part of this effort, measurements were conducted at GANIL to determine the half-life and SA branching ratio of 38Ca [8] and 30[9] using the in-flight fragment separator LISE [10]. In these experiments, the detection setup included plastic scintillators for detecting $\beta$ radiation, monitoring beam purity, and implantation onto a moving tape-drive system, as well as HPGe detectors for measuring the de-exciting $\gamma$ radiation. The tape-drive system allows separating the implantion and decay positions, which is essential to measure the decay in a background-free environment and to remove activity from long-lived beam contaminants and unstable decay products. The cycle times for these measurements were optimized for an implantation time of approximately $3t_{1/2}$ and decay times ranging from $3t_{1/2}$ for branching ratio measurements to $10t_{1/2}$ $20t_{1/2}$ for half-life measurements. The cycle time also included background measurements before implantation and after moving the activity into storage cassettes. To determine the branching ratio, a High Purity Germanium (HPGe) detector was used in coincidence with a thin plastic scintillator. The HPGe was characterized in efficiency at the 0.2% level when placed 15 cm from the decay point [11, 12]. Such a precise characterization is critical for achieving the 0.3% precision on the BR required for these measurements.
@@
A similar experiment will be conducted at GANIL in June 2025 to measure the half-life and SA BR in 18Ne [13]. This measurement will be done using the radioactive 18Ne beam produced at SPIRAL1 [14], the Isotope Separation On-Line (ISOL) facility of GANIL. The currently published branching ratio [3] for this decay has a precision of $\pm$2.7%, which is used to determine the 18Ne $ft^{0^{+}\rightarrow 0^{+}}$ value. A second measurement [15], published in a Ph.D. thesis, quotes a statistical uncertainty of ±$\pm 0.7\%$. The later was done at GANIL but the final precision on the BR was limited by statistics and affected by rate-dependent effects. The upcoming 18Ne measurement [13] will use a similar $\beta-\gamma$ detection setup as mentioned above and will be done at GANIL’s SPIRAL1 identification station [16] using a moving tape system, a plastic scintillator for $\beta$ detection and the same HPGe detector [11, 12] mentioned above. To address rate-dependent effects encountered in the previous 18Ne measurement [15], signals from the various detectors will be processed using FASTER [17], a triggerless digital data acquisition system that timestamps detected events with a precision of 2 ns.
@@
In the near future, we will conduct experiments at LISE to improve the precision on the SA branching ratio for 42Ti, 46Cr, 50Fe and 54Ni. However, beyond $Z=20$, the neutron-deficient radioactive ion beams (RIBs) that can be produced via in-flight fragmentation at LISE are not pure enough (typical purity requirement is $>99\%$) to perform the high-precision measurements required for SA $\beta$ decay studies. The purity and consequently the RIB intensity is compromised due to the overlap of momentum distribution tails from the less exotic isotones, which are produced at much higher intensities in the fragmentation process. Additionally, these $T_{Z}=-1$ parents decay to daughter nuclei that have a significant number of energetically available $J^{\pi}=1^{+}$ states. These $0^{+}\rightarrow 1^{+}$ Gamow-Teller (GT) transitions split the $\beta$ decay strength as they compete with the $0^{+}\rightarrow 0^{+}$ superallowed $\beta$ decay branch. In such cases, the branching ratio for the SA decay has to be determined indirectly by measuring all the non-analog branches and subtracting them from the full $\beta$ decay,
@@
In Eq. 2 above, $i$ runs over all non-SA $\beta$ decay branches, which need to be determined with sufficiently high precision to obtain the required $0.3\%$ precision on the SA decay branch. However, some of these non-analog branches are weaker and the de-exciting gamma radiations have high energies, as can be seen for the case for 54Ni in Fig. 4. This sets up the stage for the “Pandemonium effect” [18] when high resolution HPGe detectors are used to determine the branching ratios. This systematic effect arises from an underestimation of the $\beta$ decay strength at high energy levels, which occurs when weak $\gamma$ cascades or high-energy $\gamma$ radiations are not detected or are missed due to the low efficiency of such high resolution germanium detectors.
@@
To address these issues at LISE, the SA parent will be implanted in a segmented charged-particle detector instead of a moving tape system. To determine the half-life, a suitable time correlation will be used between the events corresponding to the implant position of the SA parent and the $\beta$ events, to avoid counting the $\beta$ radiations resulting from contaminants. To address the Pandemonium effect, gamma detectors that perform as Total Absorption Spectrometers (TAS) such as the ones being developed within the (NA)2STARS project [21] will be used to determine the branching ratio. These TAS $\gamma$ detectors are large scintillators (e.g., NaI or BaF2 ) that have high $\gamma$ detection efficiency and are used as calorimeters, covering a nearly $4\pi$ geometry to detect the entire cascade of $\gamma$ decays rather than individual $\gamma$ rays. To address contamination from radioactive daughters, the measurement cycles will be optimized for appropriate implantion and decay times with accurate forward and backwards time correlation gates between the implantation and $\beta-\gamma$ events. One such experiment is scheduled to be performed at LISE in 2026 [22] and will use a double-sided silicon strip detector (DSSSD) for implantation, plastic scintillators for $\beta$ detection, and either the DTAS detector (18 NaI(Tl) crystals) [23] or the Rocinante detector (12 BaF2 crystals) [24] as a total absorption spectrometer.
As the atomic number increases and we move further up the nuclear chart towards 98In, the SA emitters become more exotic, and based on the trend in Fig. 2, the isospin symmetry breaking corrections will increase further [6]. To test the robustness of the nuclear structure models that calculate ISB corrections, we need high-precision experimental data across the nuclear chart, especially covering large variations in the atomic number $Z$. Currently, the uncertainty in the branching ratio for the known heavier SA beta emitters, such as 66As or 70Br, is at least 10 times larger than the minimum prescribed limit [3]. As mentioned earlier, the purity of the neutron-deficient nuclei beyond 54Ni, that can be produced at LISE via in-flight fragmentation is severely limited. In addition, the refractory nature of the heavier SA emitters limits their production at SPIRAL1 via the current ISOL technique. Thus, the studies of SA emitters from 66As to 98In, focusing on the $N=Z$ cases, requires developments in RIB production and purification methods.
@@
Significant efforts are currently underway at GANIL to produce neutron-deficient RIBs at SPIRAL1, where, as mentioned earlier, the ISOL technique is being used. In this context, the on-going TULIP project [25, 26] aims to produce RIBs at SPIRAL1 via fusion-evaporation reactions, which also favors the production of neutron-deficient nuclei. As part of this project, the production of 74-78Rb+ isotopes was recently demonstrated [27] and production yields for neutron-deficient galium isotopes (including the SA emitter 62Ga), will be investigated in an upcoming beam test run [28]. Furthermore, with the beam purification techniques that will be implemented at GANIL’s upcoming low-energy facilities DESIR (Désintégration, Excitation et Stockage d’Ions Radioactifs) [29] and S3-LEB (Super Separator Spectrometer’s Low Energy Branch) [30], high-precision measurements of SA decay observables for nuclei beyond 54Ni will become possible. In this direction, S3 conducted its first commissioning run in November 2024, while DESIR is gearing up for its first commissioning run in 2027, setting up opportunities for unprecedented high-precision measurements at GANIL.
@@
The Super Separator Spectrometer of GANIL will take intense stable beams from SPIRAL2 [31, 32] and produce RIBs in-flight via the fusion evaporation method, which is favorable for producing neutron-deficient nuclei, like the SA emitters in this project. The cocktail of RIBs that is produced will be slowed down and neutralized in a gas cell [33, 34]. The neutralized atoms will be laser ionized to select and transport the RIB of interest towards the user setups. S3-LEB will also be coupled to a multi-reflection time-of-flight spectrometer called PILGRIM, which will have a mass resolution $m/\Delta m$ ranging from 1.5×$10^{5}$ to 5.2×$10^{5}$, depending on the beam emittance [35]. Studying SA $\beta$ decays at S3-LEB presents a few challenges, the resolution of which will require some developments. First, the current version of the S3-LEB gas cell has an extraction time of around 500 ms [36] which is much longer compared to the half lives of the heavier SA emitters e.g. $t_{1/2}$ for 70Br is $78.88\pm 0.31$ ms [37]. Efforts are currently being made within the S3-LEB collaboration to reduce the extraction time to $\sim 50$ ms using a fast gas cell called FRIENDS3 (Fast Radioactive Ion Extraction and Neutralization Device for S3) [38] which is being developed. The second challenge is the availability of appropriate laser ionization schemes, and laser spectroscopy studies are not done yet on 66As, 70Br, 83Tc and 90Rh [39]. The collaboration has a laser laboratory at GANIL called GISELE [40] which has the same laser systems as at S3-LEB and is coupled to an atomic beam unit that can be used to produce the stable element of interest for improving existing or developing new laser ionization schemes.

### v2

@@ 1 Introduction
Precision measurements of observables in superallowed (SA) Fermi $\beta$ transitions offer excellent probes to hunt for signatures of physics beyond the Standard Model (BSM) [1]. These decays proceed from the $J^{\pi}=0^{+}$ ground-state of the parent nucleus to the $J^{\pi}=0^{+}$ isobaric analog state in the daughter. In these decays, only the vector part of the weak interaction contributes. Consequently, the hypothesis of a conserved vector current (CVC) [2] predicts that the transition rate ($ft$ value) remains constant across all SA decays for a given isospin, irrespective of the parent-daughter pair. This hypothesis has been tested and verified to a precision of $5\times 10^{-4}$ [3] once individual $ft$ values are corrected for nuclear structure and radiative effects (denoted $\mathcal{F}t$). The validation of CVC thus allows to derive $V_{ud}$, the dominant top-row element of the Cabibbo-Kobayashi-Maskawa (CKM) matrix, that describes the relation between quark flavor eigenstates and mass eigenstates. The unitarity of this CKM matrix is fundamental to the current version of the Standard Model and is tested to the highest precision using the first-row elements. Being the largest and the most precise among the three elements, precision on $V_{ud}$ is key to performing sensitive tests of CKM unitarity. In fact, the highest precision on $V_{ud}$ currently comes from $\mathcal{F}t$ values of superallowed $0^{+}\rightarrow 0^{+}$ beta decays ($\mathcal{F}t^{0^{+}\rightarrow 0^{+}}$[3, 4]. Until recently, CKM unitarity was respected, however, a recent global re-evaluation of the $\mathcal{F}t^{0^{+}\rightarrow 0^{+}}$ values has indicated a tension with unitarity at the $2\sigma$ level [3, 4]. Deviation from unitarity raises questions on the maximal violation of parity in weak interactions, thus opening doors to possible existence of exotic scalar $(S)$ and/or tensor $(T)$ currents in the electroweak sector [3, 4]. Another effect of relaxing maximal parity violation is the possibility of right-handed interactions via the $V+A$ coupling [3]. Other scenarios that have been proposed to explain the non-unitarity of the CKM matrix, such as additional neutral gauge bosons, another generation of quarks, supersymmetry, etc, have been explored in Ref [5, 6], to quote a few.
@@
Observables in SA $0^{+}\rightarrow 0^{+}$ $\beta$ decays, namely branching ratio (BR), half-life ($t_{1/2}$), and Q-value, are used to determine the experimental $ft^{0^{+}\rightarrow 0^{+}}$ value for the decay [4]. Individual $ft^{0^{+}\rightarrow 0^{+}}$ values, shown in the left panel of Fig. 1, are corrected for nuclear medium dependent isospin symmetry breaking (ISB) and radiative effects [3] to obtain the nucleus independent $\mathcal{F}t^{0^{+}\rightarrow 0^{+}}$ values shown in the right panel of Fig. 1. The average of these $\mathcal{F}t^{0^{+}\rightarrow 0^{+}}$ values is then used to derive $V_{ud}$,
$\mathcal{F}t^{0^{+}\rightarrow 0^{+}}=ft^{0^{+}\rightarrow 0^{+}}(1+\delta^{{}^{\prime}}_{R})\left(1+\delta_{NS}-\delta_{C}\right)=\frac{K}{2V_{ud}^{2}G_{F}^{2}\left(1+\Delta_{R}\right)}.$ (1)
@@
Here, the right hand side forms a constant with $K~{}=~{}8120.2787(11)\times 10^{-10}$ GeV-4s and the nucleus-independent radiative correction $\Delta_{R}$, which is the same for all $0^{+}\rightarrow 0^{+}$ SA $\beta$ transitions [7]. In the middle term of Eq. 1, $\delta^{{}^{\prime}}_{R}$ and $\delta_{NS}$ are radiative corrections, and $\delta_{C}$ is the isospin symmetry breaking correction. $\delta^{{}^{\prime}}_{R}$ is independent of the details of the nuclear structure but depends on the atomic number Z of the daughter and the maximum energy of the beta particle emitted in the decay [4]. Computation of $\delta_{NS}$ on the contrary requires detailed description of the nuclear structure of both the initial and final nuclei, details of which can be found in Ref [3]. As the average of $\mathcal{F}t^{0^{+}\rightarrow 0^{+}}$ defines the absolute value of $V_{ud}$, it is absolutely necessary to constrain these correction terms.
[…]
@@
Currently, the $ft^{0^{+}\rightarrow 0^{+}}$ values are measured for 23 SA $\beta$ emitters, ranging from 10C $(Z=6)$ to 74Rb $(Z=37)$ [3]. However, only 15 of these, shown in Fig. 1, are used in the evaluation of $V_{ud}$ because their $ft^{0^{+}\rightarrow 0^{+}}$ values have the required precision of $0.3\%$ or better [3]. For the remaining eight cases, that are shown in the left panel of Fig. 2, the uncertainty is dominated by those on the BR and Q-value (used to determine $f$ in Fig 2).
Figure 2: Left panel: Uncertainty contribution to the corrected $\mathcal{F}t^{0^{+}\rightarrow 0^{+}}$ for the eight less precise cases of superallowed $\beta$ emitters. Figure created with data from Ref. [3]. Right panel: ISB corrections for $0^{+}\rightarrow 0^{+}$ SA decays from different models: SM-WS (2015) from Ref [3], SM-HF (1995) from Ref [8], RHF-RPA (2009) and RH-RPA (2009) from Ref [9], SV-DFT (2012) and SHZ2-DFT (2012) from Ref [10], Damgaard from Ref [12] and IVMR (2009) from Ref [13]. Figure adapted from Ref [16].
To effectively constrain the ISB corrections, a precision of $0.3\%$ or better is required on the experimental $ft$ values [3]. In Fig. 3, we demonstrate using 18Ne, how a precision of $0.3\%$ on the SA branching ratio for this decay, can help distinguish between two advanced models.
Figure 3: Top panel: $\mathcal{F}t^{0^{+}\rightarrow 0^{+}}$ obtained using two different nuclear structure models that calculate the ISB corrections for $ft^{0^{+}\rightarrow 0^{+}}$. Black data points correspond to shell model based corrections from Ref [3], referred to as SM-WS in Fig. 2. Red data points correspond to density functional based corrections of Ref [11], referred to as SV-DFT in Fig. 2. The two straight lines represent the average $\mathcal{F}t^{0^{+}\rightarrow 0^{+}}$ calculated using the 15 cases shown in Fig. 2. The inset displays the $\mathcal{F}t^{0^{+}\rightarrow 0^{+}}$ value for 18Ne determined using existing experimental data and the ISB corrections of Ref [3].
Bottom panel: ISB corrections $\delta_{C}$ taken from Ref [3] and [11]. Assuming $\delta_{C}=1$, we apply the CVC filter described in Ref [3] and propagate the uncertainties from the 18Ne $ft^{0^{+}\rightarrow 0^{+}}$ value, the average of the $\mathcal{F}t^{0^{+}\rightarrow 0^{+}}$ values for the 15 most precise SA $\beta$ emitters, as well as the theoretical uncertainty in $\delta^{{}^{\prime}}_{R}$ and $\delta_{NS}$. This yields the purple error bar, that demonstrates the potential to distinguish between the two models if the branching ratio for 18Ne can be measured with a precision of 0.3%.
@@ 2 Experimental program
At GANIL, an experimental program is currently underway aiming towards constraining ISB corrections by performing high-precision measurements of the half-life and branching ratio of the eight less-precisely known SA $\beta$ emitters. Of these eight nuclei, the first six (shown in Fig. 2) are $T_{Z}=-1$ parents, where the $\delta_{C}$ corrections are larger than those in $\beta$ decay pairs with $T_{Z}=0$ parents of the same mass. Furthermore, for the SA parent nuclei 18Ne, 30S and 42Ti - which, like 34Ar, exhibit closed neutron shell configurations - some models predict a systematically larger value of $\delta_{C}$ at shell-closures. Others, however, show no such variations (e.g., brown triangles in Fig. 2), leading to larger discrepancies between different models. Additionally, as the atomic number increases, the magnitude of $\delta_{C}$ increases, and as one can observe for the three heaviest SA emitters in Fig. 2, these corrections can range from approximately 0.1% to 1.8%. Given the sensitivity of these candidates for evaluating $\delta_{C}$, the experimental program at GANIL focuses on improving the precision of these eight known SA emitters and extending this study to the $N=Z$ SA emitters beyond 74Rb to test the $\delta_{C}$ models for larger variations in $Z$.
2.1 Past superallowed beta decay studies at GANIL
@@
As part of this effort, measurements were conducted at GANIL to determine the half-life and SA branching ratio of 38Ca [17] and 30[18] using the in-flight fragment separator LISE [19]. In these experiments, the detection setup consisted of plastic scintillators for detecting $\beta$ radiation, monitoring beam purity, and implantation onto a moving tape-drive system, as well as HPGe detectors for measuring the de-exciting $\gamma$ radiations. The tape-drive system allowed separating the implantion and decay positions, which is essential to measure the decay in a background-free environment and to remove activity from long-lived beam contaminants and unstable decay products. To determine the branching ratio, a High Purity Germanium (HPGe) detector was used in coincidence with a thin plastic scintillator. The HPGe detector was characterized in efficiency at the 0.2% level when placed 15 cm from the decay point [20, 21]. Such a precise characterization is critical for achieving the 0.3% precision on the BR required for these measurements. The incident beam intensity and cycle times for these measurements were also optimized to collect enough statistics to meet the precision requirements. The beam was implanted for $3t_{1/2}$ and decay time ranged from $3t_{1/2}$ for branching ratio measurements to $10t_{1/2}$ $20t_{1/2}$ for half-life measurements. The cycle time also included background measurements before implantation and after moving the activity into storage cassettes.
2.2 Upcoming superallowed beta decay studies at SPIRAL1 and LISE at GANIL
@@
A similar experiment will be conducted at GANIL in June 2025 to measure the half-life and SA branching ratio in 18Ne [22]. This measurement will be done using the radioactive 18Ne beam produced at SPIRAL1 (Système de Production d’Ions Radioactifs Accélérés en Ligne phase 1) [23], the Isotope Separation On-Line (ISOL) facility of GANIL. The currently published branching ratio [3] for this decay has a precision of $\pm$2.7%, which is used to determine the 18Ne $ft^{0^{+}\rightarrow 0^{+}}$ value. A second measurement [24], published in a Ph.D. thesis, quotes a statistical uncertainty of ±$\pm 0.7\%$. This second measurement was done at GANIL but the final precision on the BR was limited by statistics and affected by rate-dependent effects. The later arose from the use of peak-sensing, VXI-based analog-to-digital converters (ADCs) for measuring the branching ratio. At high counting rates, such ADCs preferentially register high-energy $\gamma$-rays, thus biasing the measured relative intensity. Furthermore, pulse pileup within the coding gate reduces the intensity derived for the strongest gamma rays. The lack of statistics also resulted due to tape breaks with the then newly commissioned moving tape system at the low-energy identification station at SPIRAL1 [25].
The upcoming 18Ne measurement [22] will use a similar $\beta-\gamma$ detection setup as mentioned in section 2.1 above, a moving tape system, a plastic scintillator for $\beta$ detection and the same HPGe detector [20, 21] mentioned above for $\gamma$ detection. To address rate-dependent effects encountered in the previous 18Ne measurement [24], signals from the various detectors will be processed using the FASTER [26] digital data acquisition (DAQ) system as opposed to the previously used VXI based analogue DAQ. FASTER is a triggerless digital DAQ that timestamps detected events with a precision of 2 ns. Furthermore, signal treament in FASTER requires less than $10\mu s$ compared to the VXI DAQ that requires several 10s of $\mu s$. This helps reduce the systematic deadtime making FASTER suitable for high-rate measurements. In addition, pulse shape analysis of the detector signals will allow to identify and accurately quantify pileup events.
@@
In the near future, we will also conduct experiments at LISE to improve the precision on the SA branching ratio for 42Ti, 46Cr, 50Fe and 54Ni. Beyond $Z=20$, the neutron-deficient radioactive ion beams (RIBs) that can be produced via in-flight fragmentation at LISE are not pure enough (typical purity requirement is $>99\%$) to perform the high-precision measurements required for SA $\beta$ decay studies. The purity and consequently the RIB intensity is compromised due to the overlap of momentum distribution tails from the less exotic isotones, which are produced at much higher intensities in the fragmentation process. Additionally, these $T_{Z}=-1$ parents decay to daughter nuclei that have a significant number of energetically available $J^{\pi}=1^{+}$ states. These $0^{+}\rightarrow 1^{+}$ Gamow-Teller (GT) transitions split the $\beta$ decay strength as they compete with the $0^{+}\rightarrow 0^{+}$ superallowed $\beta$ decay branch. In such cases, the branching ratio for the SA decay must be determined indirectly by measuring all the non-analog branches and subtracting their sum from the total $\beta$ decay,
@@
In Eq. 2 above, $i$ runs over all non-SA $\beta$ decay branches, which need to be measured with sufficiently high precision to obtain the targeted $0.3\%$ precision on the SA decay branch. Some of these non-analog branches, however, are weaker and the de-exciting gamma radiations have high energies, as can be seen for the case for 54Ni in Fig. 4. This sets up the stage for the “Pandemonium effect” [27] when high resolution HPGe detectors are used to determine the branching ratios. This systematic effect arises from an underestimation of the $\beta$ decay strength at high energy levels, which occurs when weak $\gamma$ cascades or high-energy $\gamma$ radiations are not detected or are missed due to the low efficiency of such high resolution germanium detectors.
@@
To address these issues at LISE, the SA parent will be implanted in a segmented charged-particle detector instead of a moving tape system. To determine the half-life while avoiding $\beta$ radiations from contaminants, a suitable time correlation will be applied between the events corresponding to the implant position of the SA parent and the $\beta$ decay events. To address the Pandemonium effect, gamma detectors that perform as Total Absorption Spectrometers (TAS) such as the ones being developed within the (NA)2STARS project [30] will be used to determine the branching ratio. These TAS $\gamma$ detectors are large scintillators (e.g., NaI or BaF2 ) that have high $\gamma$ detection efficiency and are used as calorimeters, covering a nearly $4\pi$ geometry to detect the entire cascade of $\gamma$ decays rather than individual $\gamma$ rays. To address contamination from radioactive daughters, the measurement cycles will be optimized for appropriate implantion and decay times with accurate forward and backwards time correlation gates between the implantation and $\beta-\gamma$ events. One such experiment is scheduled to be performed at LISE in 2026 [31] and will use a double-sided silicon strip detector (DSSSD) for implantation, plastic scintillators for $\beta$ detection, and either the DTAS detector (18 NaI(Tl) crystals) [32] or the Rocinante detector (12 BaF2 crystals) [33] as a total absorption spectrometer. For SA $\beta$ decay studies, this setup will be complemented with the previouslly mentioned HPGe detector [20, 21].
@@
As the atomic number increases and we move further up the nuclear chart towards 98In, the SA emitters become more exotic, and based on the trend in Fig. 2, the isospin symmetry breaking corrections will increase further [16]. To test the robustness of the nuclear structure models that calculate ISB corrections, we need high-precision experimental data across the nuclear chart, especially covering large variations in the atomic number $Z$. Currently, the uncertainty in the branching ratio for the known heavier SA beta emitters, such as 66As or 70Br, is at least 10 times larger than the minimum prescribed limit [3]. As mentioned earlier, the purity of the neutron-deficient nuclei beyond 54Ni, that can be produced at LISE via in-flight fragmentation is severely limited. In addition, the refractory nature of the heavier SA emitters limits their production at SPIRAL1 via the current ISOL technique. Thus, the studies of SA emitters from 66As to 98In, focusing on the $N=Z$, $T_{Z}=0$ cases, requires developments in RIB production and purification methods.
Several accelerator facilities worldwide use the ISOL technique to produce radioactive ions either by proton-induced fission and spallation (ISOLDE at CERN, SPES at INFN Legnaro, TRIUMF in Canada, ANL in USA) and light-particle induced transfer reactions (JYFL in Finland). Others produce RIBs via in-flight projectile fragmentation (GSI in Germany, NSCL in USA) similar to LISE at GANIL or in-flight fission (RIKEN in Japan) which is better suited for producing neutron-rich RIBs. These techniques are not well suited to produce the heavier neutron-deficient superallowed emitters at the required intensity and purity. As an example, 62Ga beams are available at ISOLDE and TRIUMF with intensities on the order of $10^{3}$ pps [34, 35]. For a reasonable measurement time that could be alloted at such high-demand experimental facilities, attaining a statistical precision below 1% on the SA branching will be challenging. An alternative method to produce these neutron-deficient RIBs is by fusion-evaporation reactions which has been implemented at IGISOL in Finland and will be implemented at GANIL where yields larger than $10^{4}$ pps are expected for 62Ga and 74Rb [36, 37].
@@
Significant efforts are currently underway at GANIL to produce neutron-deficient RIBs at SPIRAL1, where, as mentioned earlier, the ISOL technique is being used. In this context, the on-going TULIP project [38, 39] aims to produce RIBs at SPIRAL1 via fusion-evaporation reactions, which also favors the production of neutron-deficient nuclei. As part of this project, the production of 74-78Rb+ isotopes was recently demonstrated [40] and production yields for neutron-deficient galium isotopes (including the SA emitter 62Ga) will be investigated in an upcoming beam test run [41]. Furthermore, with the beam purification techniques that will be implemented at GANIL’s upcoming low-energy facilities DESIR (Désintégration, Excitation et Stockage d’Ions Radioactifs) [42] and S3-LEB (Super Separator Spectrometer’s Low Energy Branch) [43], high-precision measurements of SA decay observables for nuclei beyond 54Ni will become possible. In this direction, S3 conducted its first commissioning run in November 2024, while DESIR is gearing up for its first commissioning run in 2027, setting up opportunities for unprecedented high-precision measurements at GANIL.
@@
The Super Separator Spectrometer of GANIL will take intense stable beams from SPIRAL2 [44, 45] and produce RIBs in-flight via the fusion evaporation method, which is favorable for producing neutron-deficient nuclei, like the SA emitters in this project. The cocktail of RIBs that is produced will be slowed down and neutralized in a gas cell [46, 47]. The neutralized atoms will be laser-ionized to select and transport the RIB of interest towards user setups. S3-LEB will also be coupled to a multi-reflection time-of-flight spectrometer called PILGRIM, which will have a mass resolution $m/\Delta m$ ranging from $1.5\times 10^{5}$ to $5.2\times 10^{5}$, depending on the beam emittance [48]. For $10^{4}$ radioactive ions, this would allow to measure the masses with a precision of few keVs, which is at the level required for SA beta decay measurements.
Studying SA $\beta$ decays at S3-LEB, currently presents a few challenges, the resolution of which will require some developments. First, the current version of the S3-LEB gas cell has an extraction time of around 500 ms [49] which is much longer compared to the half lives of the heavier SA emitters, e.g. $t_{1/2}$ for 70Br is $78.88\pm 0.31$ ms [50]. Efforts are currently being made within the S3-LEB collaboration to reduce the extraction time to $\sim 50$ ms using a fast gas cell called FRIENDS3 (Fast Radioactive Ion Extraction and Neutralization Device for S3) [51], that is currently under development. The second challenge is the availability of appropriate laser ionization schemes, and laser spectroscopy studies are not done yet on 66As, 70Br, 83Tc and 90Rh [52]. The collaboration has a laser laboratory at GANIL called GISELE [53] which has the same laser systems as at S3-LEB and is coupled to an atomic beam unit that can be used to produce the stable element of interest for improving existing or developing new laser ionization schemes.